x+y=17 2x+4y=56
how did they get the answer 56?

Answer by stanbon(75887)

x+y=17
2x+4y=56
----
x+y = 17
x + 2y = 28
----
Subtract top from bottom and solve for "y":
y = 11
----
Solve for "x":
x + y = 17
x + 11 = 17
x = 6
--------
Cheers,
Stan H.